# Estonian WinoGrande Dataset: Comparative Analysis of LLM Performance on Human and Machine Translation

### v1

@@ 1.  Introduction
The WinoGrande dataset (sakaguchi2019winograndeadversarialwinogradschema) is a widely used benchmark for evaluating the commonsense reasoning abilities of large language models (LLMs) (team2024gemma; jiang2024mixtral; touvron2023llama; grattafiori2024llama). However, this dataset is exclusively in English, limiting evaluation of commonsense reasoning only to that language. Meanwhile, large language models have recently become increasingly multilingual, creating the need for comparable benchmarks that assess these capabilities beyond English.
@@
A common approach to creating benchmarks for other languages is to translate an existing benchmark from English to the target language. To reduce translation costs, machine translation (MT) systems are often used for this purpose  (lai-etal-2023-okapi; foroutan-etal-2023-breaking; thellmann2024multilingualllmevaluationeuropean; singh-etal-2024-aya; dang2024aya; raihan-etal-2025-mhumaneval). However, evaluation results on machine-translated data tend to differ from those on human-annotated data (kreutzer2025d). This difference is not easily predictable and may arise from multiple factors such as models exploiting translation artifacts (artetxe-etal-2020-translation) or translation inaccuracies in test data that obscure the benchmark task (plaza2024spanishllmbenchmarksmmlu). Moreover, machine translation lacks mechanisms for localization or cultural adaptation, resulting in datasets that are semantically distant from natural language use by native speakers. Finally, the WinoGrande benchmark is particularly difficult to translate across languages due to the inherent nature of the task and the strict constraints each example must satisfy.
To further examine the differences, we manually identified machine translated examples where the meaning has shifted from the original meaning. We observed that, on average, models score lower on these examples than on human-translated equivalents, indicating potential label-sentence inconsistencies introduced by machine translation.
Among the 1,767 examples in the translated WinoGrande test set, we manually corrected 89 (5% of the entire test set), while localization and adaptation was applied to 53 examples (3%). Among the remaining 1,617 examples that were not manually corrected, removed, localized or adapted, up to 15.2% lost their original semantics during machine translation. Taken together, these results demonstrate the ongoing infeasibility of using current state-of-the-art LLMs to reliably translate WinoGrande sentences into Estonian. [^1] The human translated Estonian WinoGrande dataset is available at https://huggingface.co/datasets/tartuNLP/winogrande_et.
2.  WinoGrande Dataset
The Winograd Schema Challenge (WSC) (winograd) is a pronoun reference disambiguation task proposed to evaluate commonsense knowledge in AI systems and has inspired a range of subsequent benchmarks (sakaguchi2019winograndeadversarialwinogradschema; rudinger2018genderbiascoreferenceresolution; zhang-etal-2020-winowhy) for testing AI capabilities. The original WSC is a set of few hundred expert-crafted examples intended only to assess AI capabilities. However, it was later demonstrated to be solvable with fine-tuning deep neural networks on auxiliary datasets without particular advancements in commonsense reasoning being made (kocijan-etal-2019-surprisingly; sakaguchi2019winograndeadversarialwinogradschema).
To address this issue, sakaguchi2019winograndeadversarialwinogradschema introduced WinoGrande, a large-scale ambiguity resolution dataset, in total consisting of 44k crowdsourced problems inspired by the original Winograd Schema Challenge. The dataset consists of sentences where in the first part of a sentence two nouns are mentioned and the second part contains a blank which corresponds to the mention of one of the nouns (see Table 1). Some sentences are paired (hereafter twin sentences), have a lexical overlap on 70% and share the same set of answer options (see Table 3). The objective is to decide which of the two nouns correctly fills the blank.
According to sakaguchi2019winograndeadversarialwinogradschema, the sentences are designed to meet two additional criteria. First, the answer options must be unambiguous, meaning that humans would easily be able to select the correct option without considerable effort. Secondly, the correct answer option should not be derivable solely from the local context or simple word associations. To enforce the second constraint, the authors applied a bias-reduction algorithm that extends the idea of filtering out human-detectable lexical biases to the embedding space. The algorithm identifies and removes examples that a model could solve by exploiting statistical regularities in word embeddings rather than true reasoning. Despite extensive manual and automatic filtering, the dataset has been reported to have several issues—such as instances solvable through simple word correlations, poorly written examples, sentences revealing the correct answer, or examples that are genuinely difficult to understand (kocijan2023defeatwinogradschemachallenge)—and is not completely free of artifacts (elazar-etal-2021-back).
@@ 2.  WinoGrande Dataset
Although the WinoGrande dataset has turned out to be less challenging for large language models than anticipated (Lourie_Le), it measures the progress only for English, leaving it unclear how well the success of solving ambiguity resolution tasks transfers to other languages. To answer this question, both Winograd Schema Challenge and WinoGrande have been translated to other languages (amsili-seminck-2017-google; Alhanai_2025; snaebjarnarson-etal-2022-warm). However, as amsili-seminck-2017-google have noted, translating these tasks is not a straightforward process and requires solving several linguistic challenges such as the disagreement of translated nouns in number or gender or general ambiguity of the translated schema.
Another issue with translating these datasets to other languages is the culture- and region-specific knowledge that some of these examples assume and which might be in the role of commonsense knowledge that is needed to resolve ambiguity. However, for the speakers of non-English languages, this type of knowledge can not be expected to be part of commonsense knowledge meaning that such examples must be specifically adapted for the target language. For instance, snaebjarnarson-etal-2022-warm reports having to do cultural adaptation of some of the examples. Alhanai_2025 assessed cultural appropriateness of WinoGrande translations into 11 African languages and found that 20.6% examples can be considered culturally inappropriate.
@@
This example shows the need for deliberate choices to avoid ambiguity. In English, cheap implies both low price and low quality, while its Estonian equivalent odav refers prevalently to low price. In translation, cheap is rendered as maitsetu (“tacky”) rather than literally, preserving the schema’s intended clarity and disambiguation.
@@ 3.1.  Translation Process
The translation aimed to preserve as many features of the original data set as possible. In case of twin sentences, a 70% lexical overlap was maintained between the sentences, following sakaguchi2019winograndeadversarialwinogradschema. Furthermore, since both sentences in a pair were required to share the same set of answer options, which were identical not only in lexical form but also in grammatical case, deliberate manual adjustments were necessary in the translation. Notably, achieving this is more difficult in Estonian, as its agglutinative structure complicates the preservation of morphological uniformity.
When a direct translation resulted in ambiguous schemas—often because the Estonian equivalent of a word had a broader or narrower meaning than its English counterpart—problematic words were substituted with alternatives that preserved the intended objective of the schema, similarly to amsili-seminck-2017-google. An example of this type of adjustment is presented in Table 2 (a).
Similarly to amsili-seminck-2017-google, adjustments were made in cases where a direct translation of the nouns would have resulted in disagreement in number and would have made the instance trivial to solve based on verb morphology. An example of such adjustment is presented in Table 2 (b).
@@ 3.2.  Localization, Adaptation and Error Correction
The dataset was localized by replacing geographical locations, brands, foods, activities, and animal species with culturally and regionally appropriate Estonian equivalents. A total of 53 samples were localized. Names were adapted in all samples, with English names replaced by typical Estonian names.
@@
The Cohen’s kappa between the two annotators was 0.816, and the Fleiss’ kappa between all three annotations (including the original) was 0.855, both values falling into the very high agreement range according to standard interpretation conventions (landis1977measurement). Annotator 1 found 22 items undecidable, while Annotator 2 labeled 106 tasks as such. In the majority of the cases, the annotators disagreed on which items were undecidable, with the overlap being only in eight cases. [^2] These samples are marked in our dataset.
@@ 4.  Machine Translated Dataset
We compare the human-translated dataset with two machine-translated versions produced using GPT-4o and GPT-4.1. The first version was generated with a short, generic translation prompt, while the second was produced using a longer, detailed prompt designed to address the issues observed in the first machine translated version.
@@ 4.1.  Simple Prompt
After producing the translations with the simple zero-shot translation prompt (Figure 1), we manually analyzed them for any systematic issues that could hinder the objective of the tasks or result in deviations from the required format, thereby potentially affecting the benchmark results.
The manual analysis of the machine translated data revealed the existence of translated schemas that can be solved based on grammatical cues rather than commonsense reasoning (see an example in Table 2 (b)). This can lead to predictions that are either incorrect or correct for the wrong reasons (elazar-etal-2021-back; mccoy2019right), making the schema unreliable for its intended purpose. Additionally, machine translation sometimes introduced ambiguity into the schema (see an example in Table 2 (a)).
@@
{promptbox}
Prompt 1: Simple Translation Prompt
Figure 1: Simple machine translation prompt.
@@ 4.2.  Detailed Prompt
Next, we developed two detailed prompts—one for single sentences and the other for the twin sentences—and used them to generate a revised translation of the dataset using OpenAI GPT-4.1. The prompt for single sentence items is shown in Figure 2.333The prompt for twin sentences, which is very similar with slight differences, is omitted due to space constraints. The prompts were supplemented with five single and ten twin few-shot examples, respectively. The additional samples were manually selected and translated to reflect different linguistic challenges in the task translation.
@@
{promptbox}
Prompt 2: Detailed Prompt for Single Items
Figure 2: Detailed machine translation prompt for single sentence items.
@@ 5.  Benchmarking Models
  • Five proprietary models: Gemini 2.5 Pro, Gemini 2.5 Flash gemini2025gemini25, Claude Sonnet 4.5 anthropic2025claude45sonnet, GPT-4.1 openai2024gpt4technicalreport, GPT-5 GPT-5;
  • Four open models in the moderate to large range: Llama 3.3-70B and Llama 3.1-405B grattafiori2024llama, Gemma 3-27B gemma2025gemma3, and Qwen 2.5-72B yang2025qwen3technicalreport;
  • Three open models in the smaller range: Llama 3.1-8B  grattafiori2024llama, Apertus 8B hernándezcano2025apertusdemocratizingopencompliant, EuroLLM 9B martins2024eurollmmultilinguallanguagemodels.
@@
Compared to the original English WinoGrande test set, the performance of proprietary models on the Estonian human-translated dataset was generally similar, with the average difference being only 0.4%. Performance differences were more pronounced among open models in the moderate to large model range, where on average, the models performed 5.5% better on English compared to human translated Estonian dataset. While most models performed worse on the Estonian dataset, interestingly Gemma 3-27B performed slightly better. Among smaller models, the average performance was the same on both English and Estonian datasets, with Llama 3.1-8B and Apertus-8B performed only slightly above chance level on both datasets. EuroLLM-9B was the strongest of the small models according to accuracy, and performed 1.7% better on the Estonian dataset than on the English one.
Comparing the machine translated Estonian WinoGrande versions to the human translated version reveals that all models obtain higher performance on the human translated dataset. A detailed investigation of this observation is presented in Section 6. Contrasting the results on the two machine translation versions reveals only minor differences, suggesting that the translations obtained with the detailed prompt did not yield any substantial effect on the models’ performance.
@@ 6.  Subset-Level Analysis
We analyzed the impact of cultural adaptation, and error correction in human translation on the overall results by calculating the models’ accuracies for subsets of samples that were culturally adapted (n = 53), corrected (n = 89), and the rest that were deemed semantically comparable (n = 1,617). For these analyses, we compared the results of open models of moderate to large size, as smaller models often performed near random even on the English dataset. We compared the results across three versions of the dataset—the original English, the human translated Estonian, and the machine translated Estonian created with the detailed prompt. [^4] The machine translated dataset created with the simple prompt shows very similar results.
@@ 6.1.  The Impact of Cultural Adaptation
We compared how models perform on culturally adapted examples versus those human translated examples that did not undergo such adaptation. During the human translation process, 53 schemas were thoroughly adapted by editing the content of the schema to reflect cultural relevance. We observe that, across all models, accuracies are on average 6% higher on the human translated subset that was culturally adapted (Table 5, Culturally adapted samples) compared to the human translated subset that did not undergo such adaptation (Table 5, Semantically comparable samples). Regardless of the slightly higher accuracies, we do not interpret this as evidence of the models’ stronger ability to reason based on cultural knowledge. This is because the nature of the adaptations varies: some introduce contextually relevant information—such as geographic references—that require cultural or situational understanding, while others involve changes unrelated to reasoning.

### v2

@@ 1.  Introduction
The WinoGrande dataset (Sakaguchi et al., 2019) is a widely used benchmark for evaluating the commonsense reasoning abilities of large language models (LLMs) (Gemma Team, 2024; Jiang et al., 2024; Touvron et al., 2023; Grattafiori et al., 2024). However, this dataset is exclusively in English, limiting evaluation of commonsense reasoning only to that language. Meanwhile, large language models have recently become increasingly multilingual, creating the need for comparable benchmarks that assess these capabilities beyond English.
@@
A common approach to creating benchmarks for other languages is to translate an existing benchmark from English to the target language. To reduce translation costs, machine translation (MT) systems are often used for this purpose  (Lai et al., 2023; Foroutan et al., 2023; Thellmann et al., 2024; Singh et al., 2024; Dang et al., 2024; Raihan et al., 2025). However, evaluation results on machine-translated data tend to differ from those on human-annotated data (Kreutzer et al., 2025). This difference is not easily predictable and may arise from multiple factors such as models exploiting translation artifacts (Artetxe et al., 2020) or translation inaccuracies in test data that obscure the benchmark task (Plaza et al., 2024). Moreover, machine translation lacks mechanisms for localization or cultural adaptation, resulting in datasets that are semantically distant from natural language use by native speakers. Finally, the WinoGrande benchmark is particularly difficult to translate across languages due to the inherent nature of the task and the strict constraints each example must satisfy.
@@
To further examine the differences, we manually identified machine translated examples where the meaning has shifted from the original meaning. We observe that, on average, models score lower on these examples than on human-translated equivalents, indicating potential label–sentence inconsistencies introduced by machine translation.
Among the 1,767 examples in the translated WinoGrande test set, we manually corrected 89 (5% of the entire test set), while localization and adaptation was applied to 53 examples (3%). Among the remaining semantically comparable samples, up to 15.2% lost their original semantics during machine translation. Taken together, these results demonstrate the ongoing infeasibility of using current state-of-the-art LLMs to reliably translate WinoGrande sentences into Estonian. [^1] The human translated Estonian WinoGrande dataset is available at https://huggingface.co/datasets/tartuNLP/winogrande_et
2.  WinoGrande Dataset
The Winograd Schema Challenge (WSC) (Levesque et al., 2012) is a pronoun reference disambiguation task proposed to evaluate commonsense knowledge in AI systems and has inspired a range of subsequent benchmarks (Sakaguchi et al., 2019; Rudinger et al., 2018; Zhang et al., 2020) for testing AI capabilities. The original WSC is a set of few hundred expert-crafted examples intended only to assess AI capabilities. However, it was later demonstrated to be solvable with fine-tuning deep neural networks on auxiliary datasets without particular advancements in commonsense reasoning being made (Kocijan et al., 2019; Sakaguchi et al., 2019).
To address this issue, Sakaguchi et al. (2019) introduced WinoGrande, a large-scale ambiguity resolution dataset, in total consisting of 44k crowdsourced problems inspired by the original Winograd Schema Challenge. The dataset consists of sentences where in the first part of a sentence two nouns are mentioned and the second part contains a blank which corresponds to the mention of one of the nouns (see Table 1). Some sentences are paired (hereafter twin sentences), have a lexical overlap on 70% and share the same set of answer options (see Table 5). The objective is to decide which of the two nouns correctly fills the blank.
According to Sakaguchi et al. (2019), the sentences are designed to meet two additional criteria. First, the answer options must be unambiguous, meaning that humans would easily be able to select the correct option without considerable effort. The authors report that human performance on the test set is 94%. Secondly, the correct answer option should not be derivable solely from the local context or simple word associations. To enforce the second constraint, the authors applied a bias-reduction algorithm that extends the idea of filtering out human-detectable lexical biases to the embedding space. The algorithm identifies and removes examples that a model could solve by exploiting statistical regularities in word embeddings rather than true reasoning. Despite extensive manual and automatic filtering, the dataset has been reported to have several issues—such as instances solvable through simple word correlations, poorly written examples, sentences revealing the correct answer, or examples that are genuinely difficult to understand (Kocijan et al., 2023)—and is not completely free of artifacts (Elazar et al., 2021).
WSC translations Type
French Amsili and Seminck (2017) HT
Portuguese Melo et al. (2019) HT
Japanese Shibata et al. (2015) HT
Chinese Bernard and Han (2020) HT
Hungarian Vadász and Ligeti-Nagy (2022) MT+PE
Thai Artkaew (2025) HT
Hebrew Shwartz (2024) N/A
WG translations Type
Icelandic Snæbjarnarson et al. (2022) HT
8 African languages Alhanai et al. (2025) HT
Arabic + dialects Mousi et al. (2025) MT+PE
Italian Moroni et al. (2024) MT
Kyrgys Turatali et al. (2025) MT+PE
Nepali Nyachhyon et al. (2025) MT+PE
Egyptian Shang et al. (2025) MT
Romanian Masala et al. (2024) MT
Lithuanian Nakvosas et al. (2025) MT
Korean Kim et al. (2025) HT
Basque* Corral et al. (2024) HT
Persian Farsi et al. (2025) MT+PE
Table 2: Overview of Winograd Schema Challenge (WSC) and WinoGrande (WG) dataset translations across languages. HT indicates human translation, MT indicates machine translation, and PE indicates post-editing. Asterisk (*) marks a translation of a subset of samples.
@@ 2.  WinoGrande Dataset
This example shows the need for deliberate choices to avoid ambiguity. In English, cheap implies both low price and low quality, while its Estonian equivalent odav refers mainly to low price. In translation, cheap is rendered as maitsetu (“tacky”) rather than literally, preserving the schema’s intended clarity and disambiguation.
@@
Although the WinoGrande dataset has turned out to be less challenging for large language models than anticipated (Lourie et al., 2021), it measures the progress only for English, leaving it unclear how well the success of solving ambiguity resolution tasks transfers to other languages. To answer this question, both the Winograd Schema Challenge and WinoGrande have been translated into several other languages (Table 2). Additional translations exist (Shavrina et al., 2020; Aparovich et al., 2025; Žagar and Robnik-Šikonja, 2022), for a recast version of the Winograd Schema Challenge dataset that is part of the SuperGLUE benchmark Wang et al. (2020). However, as Amsili and Seminck (2017) noted, translating these tasks is not a straightforward process and requires solving several linguistic challenges such as the disagreement of translated nouns in number or gender or general ambiguity of the translated schema.
Another issue with translating these datasets to other languages is the culture- and region-specific knowledge that some of these examples assume and which might be in the role of commonsense knowledge that is needed to resolve ambiguity. However, for the speakers of non-English languages, this type of knowledge can not be expected to be part of commonsense knowledge meaning that such examples must be specifically adapted for the target language. For instance, Snæbjarnarson et al. (2022) reports having to do cultural adaptation of some of the examples. Alhanai et al. (2025) assessed cultural appropriateness of WinoGrande translations into 11 African languages and found that 20.6% examples can be considered culturally inappropriate.
@@ 3.1.  Translation Process
The translation aimed to preserve as many features of the original dataset as possible. In case of twin sentences, a 70% lexical overlap was maintained between the sentences, following Sakaguchi et al. (2019). Furthermore, since both sentences in a pair were required to share the same set of answer options, which were identical not only in lexical form but also in grammatical case, deliberate manual adjustments were necessary in the translation. Notably, achieving this is more difficult in Estonian, as its agglutinative structure complicates the preservation of morphological uniformity.
When a direct translation resulted in ambiguous schemas—often because the Estonian equivalent of a word had a broader or narrower meaning than its English counterpart—problematic words were substituted with alternatives that preserved the intended objective of the schema, similarly to Amsili and Seminck (2017). An example of this type of adjustment is presented in Table 3 (A).
Similarly to Amsili and Seminck (2017), adjustments were made in cases where a direct translation of the nouns would have resulted in disagreement in number and would have made the instance trivial to solve based on verb morphology. An example of such adjustment is presented in Table 3 (B).
@@ 3.2.  Localization, Adaptation and Error Correction
The dataset was localized by replacing geographical locations, brands, foods, activities, and animal species with culturally and regionally appropriate Estonian equivalents. A total of 53 samples were localized. As a result of the localization process, two types of samples can be identified: those in which the culturally adapted information is necessary for reasoning ( Table 4, A) , and those in which it is not ( Table 4, B). Names were adapted in all samples, with English names replaced by typical Estonian names.
English Schema Culturally Adapted Estonian Schema
(A) Samantha lived in the city while Patricia lived in the desert, so _ thought it common to see a cactus in their yard.
Option 1: Samantha
Option 2: Patricia
Sandra elas mandril, kuid Piret elas saarel. Seega oli _ jaoks oma aias kadaka nägemine tavapärane.
Option 1: Sandra
Option 2: Piret
(B) Laura beat Erin in a game of Mortal Kombat, then _ said congratulations on the win.
Option 1: Laura
Option 2: Erin
Laura tegi Erele lauamängus „Mees, kes teadis ussisõnu“ pähe, mispeale õnnitles _ teda võidu puhul.
Option 1: Laura
Option 2: Ere
Table 4: Examples of culturally adapted WinoGrande schemas in English and Estonian. In Sample A, correct resolution requires cultural knowledge: the original contrasts city and desert (e.g., cacti in deserts), while the Estonian version replaces these with mainland (mandril) and island (saarel), and cactus with juniper (kadaka), a plant associated with Estonian islands. In Sample B, the game name is culturally adapted, but this knowledge is not necessary for reasoning.
@@
The Cohen’s kappa between the two annotators is 0.816, and the Fleiss’ kappa between all three annotations (including the original) was 0.855, both values falling into the very high agreement range according to standard interpretation conventions (Landis and Koch, 1977). The accuracy was 95% for Annotator 1 and 92% for Annotator 2. Annotator 1 found 22 items undecidable, while Annotator 2 labeled 106 tasks as such. In the majority of the cases, the annotators disagreed on which items were undecidable, with the overlap being only in eight cases. [^2] These samples are marked in our dataset.
@@ 4.  Machine Translated Dataset
We compare the human-translated dataset with two machine-translated versions produced using GPT-4o and GPT-4.1. The first version was generated with a short, generic translation prompt, while the second was produced using longer, detailed prompts designed to address the issues observed in the first machine translated version.
@@ 4.1.  Simple Prompt
After producing the translations with the simple zero-shot translation prompt (A.1), we manually analyzed them for any systematic issues that could hinder the objective of the tasks or result in deviations from the required format, thereby potentially affecting the benchmark results.
The manual analysis of the machine translated data revealed the existence of translated schemas that can be solved based on grammatical cues rather than commonsense reasoning (see an example in Table 3 (B)). This can lead to predictions that are either incorrect or correct for the wrong reasons (Elazar et al., 2021; McCoy et al., 2019), making the schema unreliable for its intended purpose. Additionally, machine translation sometimes introduced ambiguity into the schema (see an example in Table 3 (A)).
@@ 4.2.  Detailed Prompt
Next, we developed two detailed prompts—one for single sentences (A.2) and the other for the twin sentences (A.3)—and used them to generate a revised translation of the dataset using OpenAI GPT-4.1. The prompts were supplemented with five single and ten twin few-shot examples, respectively. The additional samples were manually selected and translated to reflect different linguistic challenges in the task translation.
@@ 5.  Benchmarking Models
  • Five proprietary models: Gemini 2.5 Pro, Gemini 2.5 Flash Gemini Team, Google (2025), Claude Sonnet 4.5 Anthropic (2025), GPT-4.1 OpenAI et al. (2024), GPT-5 OpenAI (2025);
  • Four open models in the moderate to large range: Llama 3.3-70B and Llama 3.1-405B Llama Team (2024), Gemma 3-27B Gemma Team (2025), and Qwen 2.5-72B Yang et al. (2025);
  • Three open models in the smaller range: Llama 3.1-8B  Llama Team (2024), Apertus 8B Hernández-Cano et al. (2025), EuroLLM 9B Martins et al. (2024).
@@
Compared to the original English WinoGrande test set, the performance of proprietary models on the Estonian human-translated dataset was generally similar, with the average difference being only 0.4%.
Performance differences were more pronounced among open models in the moderate to large model range, where on average, the models performed 5.5% better on English compared to human translated Estonian dataset. While most models performed worse on the Estonian dataset, interestingly Gemma 3-27B performed slightly better.
Among smaller models, the average performance was the same on both English and Estonian datasets, with Llama 3.1-8B and Apertus-8B performing only slightly above chance level on both datasets. EuroLLM-9B was the strongest of the small models according to accuracy, and performed 1.7% better on the Estonian dataset than on the English one.
Comparing the machine translated Estonian WinoGrande versions to the human translated version reveals that all models obtain higher performance on the human translated dataset. A detailed investigation of this observation is presented in Section 6.
Contrasting the results on the two machine translation versions reveals only minor differences, suggesting that the translations obtained with the detailed prompt did not yield an substantial effect on the models’ performance.
@@ 6.  Subset-Level Analysis
We analyzed the impact of cultural adaptation, and error correction in human translation on the overall results by calculating the models’ accuracies for subsets of samples that were culturally adapted (n = 53), corrected (n = 89), and the rest that were deemed semantically comparable (n = 1,617). The 8 samples that both annotators found ambiguous were excluded from the subset analysis. For these analyses, we compared the results of open models of moderate to large size, as smaller models often performed near random even on the English dataset. We compared the results across three versions of the dataset—the original English, the human translated Estonian, and the machine translated Estonian created with the detailed prompt. [^3] The machine translated dataset created with the simple prompt shows very similar results.
@@ 6.1.  The Impact of Cultural Adaptation
We compared how models perform on culturally adapted examples versus those human translated examples that did not undergo such adaptation. During the human translation process, 53 schemas were thoroughly adapted by editing the content of the schema to reflect cultural relevance. We observe that, across all models, accuracies are on average 6% higher on the human translated subset that was culturally adapted (Table 7, Culturally adapted samples) compared to the human translated subset that did not undergo such adaptation (Table 7, Semantically comparable samples). Regardless of the slightly higher accuracies, we do not interpret this as evidence of the models’ stronger ability to reason based on cultural knowledge. This is because the nature of the adaptations varies: some introduce contextually relevant information—such as geographic references—that require cultural or situational understanding, while others involve changes unrelated to reasoning (Table 4).